OpenStudy (anonymous):

Dione Industries bought a scanner for $800. It is expected to depreciate at a rate of 10% per year. What will the value of the scanner be in 2 years? Answer $648 $720 $889 $968

OpenStudy (anonymous):

648

OpenStudy (anonymous):

did u use your calculator

OpenStudy (anonymous):

First year it depreciates by 10% i. 10/100 * 800 = 80 its value after the first year is 800 -80 = 720 2nd year it depreciates also by 10% i.e. 10/100 * 720 = 72 So its value will be 720 -72 = 648

OpenStudy (anonymous):

No I just calculated manually

OpenStudy (anonymous):

a minor correction: If the number of years are too large you can create a formula easily as follows: Let the original value be Y Let x% be the percentage value after it depreciates. then (x/100)^n is the percentage of the remaining true value after depreciation in n years. Then its true value after n years is given by the following formula y*(x/100)^n In this case x=90 ( i.e it depreciates by 10% so the remaining percentage value is 90%) x=90 n=2 y=800 so after 2 years its true value will be given by 800*(x/100)^2 = 800*(90*90)/(100*100) = 9*9*8 = 648

OpenStudy (anonymous):

Nothing wrong with the answer above, and it may be the answer anticipated by the author of the question. Sometimes, these problems are modeled by an exponential model\[V(t) =V_0 e ^{-rt}\]If you substitute for values in this problem,\[V(2)=800e ^{-.20}\approx655\]

OpenStudy (anonymous):

thanks :)